# Towards Privacy-Guaranteed Label Unlearning in Vertical Federated Learning: Few-Shot Forgetting without Disclosure

### v2

@@ Abstract
This paper addresses the critical challenge of unlearning in Vertical Federated Learning (VFL), a setting that has received far less attention than its horizontal counterpart. Specifically, we propose the first method tailored to label unlearning in VFL, where labels play a dual role as both essential inputs and sensitive information. To this end, we employ a representation-level manifold mixup mechanism to generate synthetic embeddings for both unlearned and retained samples. This is to provide richer signals for the subsequent gradient-based label forgetting and recovery steps. These augmented embeddings are then subjected to gradient-based label forgetting, effectively removing the associated label information from the model. To recover performance on the retained data, we introduce a recovery-phase optimization step that refines the remaining embeddings. This design achieves effective label unlearning while maintaining computational efficiency. We validate our method through extensive experiments on diverse datasets, including MNIST, CIFAR-10, CIFAR-100, ModelNet, Brain Tumor MRI, COVID-19 Radiography, and Yahoo Answers demonstrate strong efficacy and scalability. Overall, this work establishes a new direction for unlearning in VFL, showing that re-imagining mixup as an efficient mechanism can unlock practical and utility-preserving unlearning. The code is publicly available at https://github.com/bryanhx/Towards-Privacy-Guaranteed-Label-Unlearning-in-Vertical-Federated-Learning
@@ 1 Introduction
However, the need for synchronous processing stands out as a paramount constraint in VFL architectures. In VFL, different parties possess unique feature sets over a common sample ID space, necessitating coordinated training steps in which intermediate results are exchanged and aligned at precise points. This rigorous coordination implies that all participating models must wait for the slowest participant to complete its step before proceeding to the next round. Consequently, this synchronization requirement significantly compounds the difficulties associated with unlearning efficiency in the context of Vertical Federated Unlearning (VFU).
@@
Our proposed unlearning solution offers three key advantages: (i) Performance-Preserving Unlearning: It requires only a small subset of labeled public data, significantly minimizing the degradation of the retain classes performance; (ii) Enhanced Unlearning Effectiveness: By leveraging the manifold mixup technique, it achieves effective unlearning with minimal data; and (iii) Computational Efficiency: The proposed unlearning process is highly efficient, completing within seconds. The primary contributions of this work are as follows:
@@
    To the best knowledge, this is the first work to address the unlearning of labels in VFL.
@@
    We propose a novel few-shot label-unlearning method that removes labels from both active and passive models in VFL using only a limited amount of public data. With this, our approach enhances unlearning effectiveness through the representation-level manifold mixup mechanism, ensuring a more robust and efficient unlearning process (see Section  4).
@@ 2.1 Machine Unlearning (MU)
MU was first proposed by (Cao & Yang, 2015) to remove specific data from a model without full retraining (Garg et al., 2020; Chen et al., 2021). It includes exact and approximate unlearning. Exact methods like SISA (Bourtoule et al., 2021) and ARCANE (Yan et al., 2022) partition data, train sub-models per partition, and retrain only the affected sections during unlearning. Approximate unlearning includes methods like fine-tuning (Golatkar et al., 2020a; Jia et al., 2023), random label updates (Graves et al., 2021; Chen et al., 2023), noise injection (Tarun et al., 2024; Huang et al., 2021), gradient ascent (Goel et al., 2022; Choi & Na, 2023; Abbasi et al., 2023; Hoang et al., 2024), knowledge distillation (Chundawat et al., 2023b; Zhang et al., 2023b; Kurmanji et al., 2023), and weight scrubbing (Golatkar et al., 2020a; b; 2021; Guo et al., 2020; Foster et al., 2024a). While these rely on data during unlearning, Chundawat et al. (Chundawat et al., 2023a) propose zero-shot unlearning, and Yoon et al. (Yoon et al., 2022) introduce a few-shot approach using model inversion.
@@ 2.2 Federated Unlearning (FU)
In FU, most of the existing works are focusing in the horizontal environment (Wu et al., 2022; Gu et al., 2024a; Zhao et al., 2023; Romandini et al., 2024; Liu et al., 2024b; Zhang et al., 2023a; Su & Li, 2023; Ye et al., 2024; Gao et al., 2024; Cao et al., 2023b; Yuan et al., 2023; Alam et al., 2024; Li et al., 2023; Halimi et al., 2022; Xia et al., 2023; Wang et al., 2023; Dhasade et al., 2023; Liu et al., 2022; Zhao et al., 2024; Wang et al., 2022; Gu et al., 2024b; Liu et al., 2021; Cao et al., 2023a). The research on horizontal federated unlearning (HFU) mainly focuses on label (Wang et al., 2022; Zhao et al., 2024), client (Liu et al., 2021; Yuan et al., 2023; Zhang et al., 2023a; Gao et al., 2024; Ye et al., 2024; Su & Li, 2023; Cao et al., 2023a; Wu et al., 2022) and sample unlearning (Liu et al., 2022; Dhasade et al., 2023). While HFU has been widely studied, vertical federated unlearning (VFU) remains underexplored. Existing works mainly focus on passive party unlearning, such as logistic regression (Deng et al., 2023), gradient boosting (Li et al., 2024), and deep learning models with fast retraining (Wang et al., 2024) or backdoor defence (Han et al., 2025). However, little attention has been given to active-party-initiated unlearning with all parties actively engaged in VFL.
@@ 3.1 General Setup
VFL training. Vertical Federated Learning allows parties with different user attributes to collaboratively train a model without sharing raw data (Yang et al., 2019; Liu et al., 2024a). The active party holds the labels and the active model, while passive parties provide features and passive models, enabling performance gains while preserving privacy. We assume that a VFL setting consists of one active party $P_{0}$ and $K$ passive parties $\{P_{1},\cdots,P_{K}\}$ who collaboratively train a VFL model $\Theta=(\theta_{1},\cdots,\theta_{K},\omega)$ to optimize:
@@
in which passive party $P_{k}$ owns features $x_{k}=(x_{k,1},\cdots,x_{k,n})$ and the passive model $G_{\theta_{k}}$, the active party owns the labels $y=\{y_{1},\cdots,y_{m}\}$ and active model $F_{\omega}$. Before training, parties use Private Set Intersection (PSI) to align data records by the same ID. During training, each passive party $k$ computes a forward embedding $H_{k}$ from its features and sends it to the active party. The active party concatenates embeddings $\{H_{k}\}_{k=1}^{K}$ into $H$, generates outputs, computes the loss, and updates its model using $\frac{\partial\textit{L}}{\partial\omega}$. It then sends $\frac{\partial\ell}{\partial H_{k}}$ to passive parties, which compute $\frac{\partial\ell}{\partial\theta_{k}}=\frac{\partial\ell}{\partial H_{k}}\cdot\frac{\partial H_{k}}{\partial\theta_{k}}$ to update their models. Please refer to Appendix A.2 for table of notations.
Label unlearning in VFL. When the active party requests to unlearn some sensitive labels $y^{u}$, where the corresponding unlearn feature is $\{x^{u}_{k}\}_{k=1}^{K}:=\{\{x_{k,i}^{u}\}_{i=1}^{n_{u}}\}_{k=1}^{K}$. The active party aims to remove the influence of $y^{u}$ on both the active model $F_{\omega}$ and $K$ passive models $\{G_{\theta_{k}}\}_{k=1}^{K}$.
Label unlearning in VFL refers to the process of efficiently and securely removing label information from a VFL system. Specifically, the unlearned passive model of passive party $k$, denoted as $\theta^{u}_{k}$, and the unlearned active model, denoted as $\omega^{u}$, are obtained through the application of an unlearning mechanism $\mathcal{U}$, as follows:
$\theta^{u}_{k}=\mathcal{U}(\theta_{k},y^{u}),\quad\omega^{u}=\mathcal{U}(\omega,y^{u}),$
where $\theta_{k}$ and $\omega$ represent the passive model of passive party $k$ and active model before unlearning, respectively, and $\mathbf{g}^{u}$ are the gradients associated with the unlearned label $y^{u}$.
Building on (Bourtoule et al., 2021), label unlearning in VFL should meet three goals: i) Performance-Preserving Unlearning: remove the impact of specific labels without affecting other data; ii) Enhanced Unlearning Effectiveness: achieve this with minimal data; and iii) Computational Efficiency: avoid full model retraining.
@@ 4 Few-Shot Label Unlearning
This section outlines the proposed few-shot label unlearning method, as shown in Figure 1 and Algorithm 1. The approach involves three steps: (1) augmenting the forward embedding with manifold mixup to address limited labeled data (Section 4.1); (2) using gradient ascent on the augmented embedding to guide label removal in both passive and active models (Section 4.2); and (3) improving model accuracy on the retained labels (Section 4.3).
@@ 4.1 Vertical Manifold Mixup
By working with fewer samples, the steps involved in the unlearning algorithm, such as forward embeddings execution and gradient updates, can be executed much faster. However, this small labeled dataset with unlearned label, denoted as $\mathcal{D}^{p,u}=\{(x_{1,i}^{p,u},y^{p,u}_{i}),\cdots,(x_{K,i}^{p,u},y^{p,u}_{i})\}_{i=1}^{n_{p,u}}$, is insufficient for an effective unlearning (see Fig 3). Consequently, this scenario can be framed as a few-shot unlearning problem, in which a minimal set of labels is used to unlearn all associated labels. Please refer to Appen. A.4 for an explanation of why a small number of samples (up to 40 per label, as shown in Tab. 9 and 10) is sufficient to achieve performance comparable to using the full dataset.
@@
$\begin{split}&\max_{\omega,\theta_{1},\cdots,\theta_{K}}\frac{1}{{n_{p,u}}^{2}}\sum_{i,j=1}^{n_{p,u}}\ell(F_{\omega}\circ(\text{Mix}_{\lambda}(G_{\theta_{1}}(x_{1,i}^{p,u}),G_{\theta_{1}}(x_{1,j}^{p,u})),\\ &\quad\cdots,\text{Mix}_{\lambda}(G_{\theta_{K}}(x_{K,i}^{p,u}),G_{\theta_{K}}(x_{K,j}^{p,u})),\text{Mix}_{\lambda}(y_{i}^{p,u},y_{j}^{p,u})),\end{split}$
@@
The mixed coefficient $\lambda$ ranges from 0 to 1. The advantage of the manifold mixup approach lies in its ability to flatten the state distributions (Verma et al., 2019). Once each passive party $k$ generates its local embeddings $\{H^{p,u}_{k}=G_{\theta}(x_{k}^{p,u})\}$, these embeddings are transmitted to the active party. The active party then constructs a set of synthesized embeddings $\vec{H}_{k}^{u}$ by performing manifold mixup among embeddings originating from the same passive party, ensuring consistency in the $\lambda$ value while avoiding any coordination among passive parties. Similarly, for the small data with the remaining label $\mathcal{D}^{p,r}$, we also implement the manifold mixup to obtain the $\vec{H}_{k}^{r}$ and corresponding label $\vec{y}^{r}$ for the remaining accuracy recovery step.
@@ 4.2 Vertical Gradient-based Label Unlearning
$\min_{\omega}\ell(F_{\omega}(\vec{H}^{u}),\vec{y}^{u})=\ell(F_{\omega}([\vec{H}_{1},\cdots,\vec{H}_{K}^{u}]),\vec{y}^{u}),$ (3)
@@
Unlearning for passive model $G_{\omega_{k}}$. Subsequently, the active party computes the gradients $\frac{\partial\ell}{\partial\vec{H}_{k}^{u}}$ in accordance with Equation 3 and transmits these gradients to the corresponding passive party $k$. Finally, the passive party $k$ updates the passive model $G_{\theta_{k}}$ using the following expression:
@@ 5.1 Experiment Setup
Datasets & Models. We conduct experiments on seven datasets: MNIST (Lecun et al., 1998), CIFAR10, CIFAR100 (Krizhevsky et al., 2009), ModelNet (Wu et al., 2015), Brain Tumor MRI (Wang et al., 2024), COVID-19 Radiography (Rahman, 2022) and Yahoo Answers dataset (Fu et al., 2022a). We adopt ResNet18 on the datasets MNIST, CIFAR10, CIFAR100, ModelNet, Brain Tumor MRI and COVID-19 Radiography. We adopt MixText (Chen et al., 2020) on the Yahoo Answers dataset. Additionally, we extend our experiments with Vgg16 on the dataset CIFAR10 and CIFAR100. Experiments are repeated over five random trials, and results are reported as mean and standard deviation. We conduct experiments on a single NVIDIA A100 GPU.
@@ 2. $i$
Evaluation for different privacy-preserving VFL methods: Our method remains effective when applied to privacy-preserving VFL models. To validate this, we evaluate performance under Differential Privacy and Gradient Compression, two widely adopted mechanisms in VFL training (Fu et al., 2022b). The results confirm that our approach maintains robustness, underscoring its practicality for real-world deployments. Please refer to Appendix A.5 for a complete result.

### v3

@@ Abstract
This paper addresses the critical challenge of unlearning in Vertical Federated Learning (VFL), a setting that has received far less attention than its horizontal counterpart. Specifically, we propose the first method tailored to label unlearning in VFL, where labels play a dual role as both essential inputs and sensitive information. To this end, we employ a representation-level manifold mixup mechanism to generate synthetic embeddings for both unlearned and retained samples. This is to provide richer signals for the subsequent gradient-based label forgetting and recovery steps. These augmented embeddings are then subjected to gradient-based label forgetting, effectively removing the associated label information from the model. To recover performance on the retained data, we introduce a recovery-phase optimization step that refines the remaining embeddings. This design achieves effective label unlearning while maintaining computational efficiency. We validate our method through extensive experiments on diverse datasets, including MNIST, CIFAR-10, CIFAR-100, ModelNet, Brain Tumor MRI, COVID-19 Radiography, and Yahoo Answers demonstrate strong efficacy and scalability. Overall, this work establishes a new direction for unlearning in VFL, showing that re-imagining mixup as an efficient mechanism can unlock practical and utility-preserving unlearning. The code is publicly available at https://github.com/bryanhx/Towards-Privacy-Guaranteed-Label-Unlearning-in-Vertical-Federated-Learning.
@@ 1 Introduction
However, the need for synchronous processing is a paramount constraint in VFL architectures. In VFL, different parties possess unique feature sets over a common sample ID space, necessitating coordinated training steps in which intermediate results are exchanged and aligned at precise points. This rigorous coordination implies that all participating models must wait for the slowest participant to complete its step before proceeding to the next round. Consequently, this synchronization requirement significantly compounds the difficulties associated with unlearning efficiency in the context of Vertical Federated Unlearning (VFU).
@@
Our proposed unlearning solution offers three key advantages: (i) Performance-Preserving Unlearning: It requires only a small subset of labeled public data, significantly minimizing the degradation of the retain classes performance; (ii) Enhanced Unlearning Effectiveness: By leveraging the manifold mixup technique, it achieves effective unlearning with minimal data; and (iii) Computational Efficiency: The proposed unlearning process is highly efficient, completing within seconds.
The primary contributions of this work are as follows:
@@
    To the best knowledge, this is the first work to address label unlearning in VFL.
@@
    We propose a novel few-shot label unlearning method that removes labels from both active and passive models in VFL using only a limited amount of public data. As a result, our approach enhances unlearning effectiveness through a representation-level manifold mixup mechanism, ensuring a more robust and efficient unlearning process (see Section  4).
@@ 2.1 Machine Unlearning (MU)
MU was first proposed by (Cao and Yang, 2015) to remove specific data from a model without full retraining (Garg et al., 2020; Chen et al., 2021). It includes exact and approximate unlearning. Exact methods like SISA (Bourtoule et al., 2021) and ARCANE (Yan et al., 2022) partition data, train sub-models per partition, and retrain only the affected sections during unlearning. Approximate unlearning includes methods like fine-tuning (Golatkar et al., 2020a; Jia et al., 2023), random label updates (Graves et al., 2021; Chen et al., 2023), noise injection (Tarun et al., 2024; Huang et al., 2021), gradient ascent (Goel et al., 2022; Choi and Na, 2023; Abbasi et al., 2023; Hoang et al., 2024), knowledge distillation (Chundawat et al., 2023a; Zhang et al., 2023b; Kurmanji et al., 2023), and weight scrubbing (Golatkar et al., 2020a; b; 2021; Guo et al., 2020; Foster et al., 2024a). While these rely on data during unlearning, Chundawat et al. (Chundawat et al., 2023b) propose zero-shot unlearning, and Yoon et al. (Yoon et al., 2022) introduce a few-shot approach using model inversion.
@@ 2.2 Federated Unlearning (FU)
In FU, most of the existing works are focusing in the horizontal environment (Wu et al., 2022; Gu et al., 2024a; Zhao et al., 2023; Romandini et al., 2024; Liu et al., 2024b; Zhang et al., 2023a; Su and Li, 2023; Ye et al., 2024; Gao et al., 2024; Cao et al., 2023b; Yuan et al., 2023; Alam et al., 2024; Li et al., 2023; Halimi et al., 2022; Xia et al., 2023; Wang et al., 2023; Dhasade et al., 2023; Liu et al., 2022; Zhao et al., 2024; Wang et al., 2022; Gu et al., 2024b; Liu et al., 2021; Cao et al., 2023a). The research on horizontal federated unlearning (HFU) mainly focuses on label (Wang et al., 2022; Zhao et al., 2024), client (Liu et al., 2021; Yuan et al., 2023; Zhang et al., 2023a; Gao et al., 2024; Ye et al., 2024; Su and Li, 2023; Cao et al., 2023a; Wu et al., 2022) and sample unlearning (Liu et al., 2022; Dhasade et al., 2023). While HFU has been widely studied, vertical federated unlearning (VFU) remains underexplored. Existing works mainly focus on passive party unlearning, such as logistic regression (Deng et al., 2023), gradient boosting (Li et al., 2024), and deep learning models with fast retraining (Wang et al., 2024) or backdoor defence (Han et al., 2025). However, little attention has been given to active-party-initiated unlearning with all parties actively engaged in VFL.
@@ 3.1 General Setup
VFL training. Vertical Federated Learning (VFL) allows parties with different user attributes to collaboratively train a model without sharing raw data (Yang et al., 2019; Liu et al., 2024a). The active party holds the labels and the active model, while passive parties provide features and passive models, enabling performance gains while preserving privacy. We assume that a VFL setting consists of one active party $P_{0}$ and $K$ passive parties $\{P_{1},\cdots,P_{K}\}$ who collaboratively train a VFL model $\Theta=(\theta_{1},\cdots,\theta_{K},\omega)$ to optimize:
@@
We denote the empirical training objective in Equation 1 as
$\mathcal{L}(\omega,\theta_{1},\ldots,\theta_{K})=\frac{1}{n}\sum_{i=1}^{n}\ell(F_{\omega}\circ(G_{\theta_{1}}(x_{1,i}),\ldots,G_{\theta_{K}}(x_{K,i})),y_{i}).$
where passive party $P_{k}$ owns features $x_{k}=(x_{k,1},\cdots,x_{k,n})$ and the passive model $G_{\theta_{k}}$; while the active party owns the labels $y=\{y_{1},\cdots,y_{n}\}$ and active model $F_{\omega}$. Before training, both parties use Private Set Intersection (PSI) to align data records by the same ID. During training, each passive party $k$ computes a forward embedding $H_{k}$ from its features and sends it to the active party. The active party concatenates embeddings $\{H_{k}\}_{k=1}^{K}$ into $H$, generates outputs, computes $\mathcal{L}$, and updates its model using $\frac{\partial\mathcal{L}}{\partial\omega}$. It then sends $\frac{\partial\mathcal{L}}{\partial H_{k}}$ to passive parties, which compute $\frac{\partial\mathcal{L}}{\partial\theta_{k}}=\frac{\partial\mathcal{L}}{\partial H_{k}}\cdot\frac{\partial H_{k}}{\partial\theta_{k}}$ to update their models. Please refer to Appendix A.2 for the table of notations.
Label unlearning in VFL. When the active party requests unlearning for a subset of samples with indices $\mathcal{I}^{u}\subseteq\{1,\dots,n\}$, we define the unlearned dataset as $\mathcal{D}^{u}=\{(x_{1,i},\dots,x_{K,i},y_{i})\}_{i\in\mathcal{I}^{u}}$. For each passive party $P_{k}$, the corresponding feature subset is $x_{k}^{u}=(x_{k,i})_{i\in\mathcal{I}^{u}}$.
The objective of label unlearning is to efficiently remove the influence of $\mathcal{D}^{u}$ from the trained VFL system without requiring full retraining. Formally, given an unlearning mechanism $\mathcal{U}$ acting on the trained model $\Theta=(\theta_{1},\dots,\theta_{K},\omega)$, the updated parameters are
$(\theta_{1}^{u},\dots,\theta_{K}^{u},\omega^{u})=\mathcal{U}(\Theta;\mathcal{D}^{u}).$
Following standard machine unlearning principles (Bourtoule et al., 2021), an effective unlearning method should satisfy: i) Performance-Preserving Unlearning, removing the impact of $\mathcal{D}^{u}$ without degrading performance on retained data; ii) Enhanced Unlearning Effectiveness, requiring minimal auxiliary data; and iii) Computational Efficiency, avoiding full model retraining.
@@ 4 Few-Shot Label Unlearning
This section outlines our proposed few-shot label unlearning method, as shown in Figure 1 and Algorithm 1. The approach involves three steps: (1) augmenting the forward embedding with manifold mixup to address limited labeled data (Section 4.1); (2) using gradient ascent on the augmented embedding to guide label removal in both passive and active models (Section 4.2); and (3) improving model accuracy on the retained labels (Section 4.3).
@@ 4.1 Vertical Manifold Mixup
With fewer samples, the steps involved in the unlearning algorithm, such as forward embeddings execution and gradient updates, can be executed much faster. However, this small labeled unlearning set, denoted as $\mathcal{D}^{p,u}=\{(\{x^{p,u}_{k,i}\}_{k=1}^{K},\,y^{p,u}_{i})\}_{i=1}^{n_{p,u}}$ is often insufficient for effective unlearning (see Figure 3). Consequently, we cast this setting as a few-shot unlearning problem, where only a small number of labeled unlearning samples are available to remove the corresponding label information from the model. Please refer to Appendix A.4 for an explanation of why a small number of samples (up to 40 per label, as shown in Tables 9 and 10) is sufficient to achieve performance comparable to using the full dataset.
@@
$\begin{split}&\max_{\omega,\theta_{1},\ldots,\theta_{K}}\frac{1}{n_{p,u}^{2}}\sum_{i,j=1}^{n_{p,u}}\ell\!\left(F_{\omega}\!\left(\text{Mix}_{\lambda}(G_{\theta_{1}}(x^{p,u}_{1,i}),G_{\theta_{1}}(x^{p,u}_{1,j})),\ldots,\text{Mix}_{\lambda}(G_{\theta_{K}}(x^{p,u}_{K,i}),G_{\theta_{K}}(x^{p,u}_{K,j}))\right),\text{Mix}_{\lambda}(y^{p,u}_{i},y^{p,u}_{j})\right).\end{split}$
@@
The mixed coefficient $\lambda$ ranges from 0 to 1. The advantage of the manifold mixup approach lies in its ability to flatten the state distributions (Verma et al., 2019). Once each passive party $k$ generates its local embeddings $H^{p,u}_{k,i}=G_{\theta_{k}}(x^{p,u}_{k,i})$, these embeddings are transmitted to the active party. The active party then constructs a set of synthesized embeddings $\vec{H}_{k}^{u}$ by performing manifold mixup among embeddings originating from the same passive party, ensuring consistency in the $\lambda$ value while avoiding any coordination among passive parties. Similarly, for the small data with the remaining label $\mathcal{D}^{p,r}$, we also implement the manifold mixup to obtain the $\vec{H}_{k}^{r}$ and corresponding label $\vec{y}^{r}$ for the remaining accuracy recovery step.
@@ 4.2 Vertical Gradient-based Label Unlearning
$\max_{\omega}\ell(F_{\omega}(\vec{H}^{u}),\vec{y}^{u})=\ell(F_{\omega}([\vec{H}_{1},\cdots,\vec{H}_{K}^{u}]),\vec{y}^{u}),$ (3)
@@
Unlearning for passive model $G_{\theta_{k}}$. Subsequently, the active party computes the gradients $\frac{\partial\ell}{\partial\vec{H}_{k}^{u}}$ in accordance with Equation 3 and transmits these gradients to the corresponding passive party $k$. Finally, the passive party $k$ updates the passive model $G_{\theta_{k}}$ using the following expression:
@@ 5.1 Experiment Setup
Datasets & Models. We conduct experiments on seven datasets: MNIST (Lecun et al., 1998), CIFAR10, CIFAR100 (Krizhevsky et al., 2009), ModelNet (Wu et al., 2015), Brain Tumor MRI (Wang et al., 2024), COVID-19 Radiography (Rahman, 2022) and Yahoo Answers dataset (Fu et al., 2022b). We adopt ResNet18 on the datasets MNIST, CIFAR10, CIFAR100, ModelNet, Brain Tumor MRI and COVID-19 Radiography. We adopt MixText (Chen et al., 2020) on the Yahoo Answers dataset. Additionally, we extend our experiments with Vgg16 on the dataset CIFAR10 and CIFAR100. Experiments are repeated over five random trials, and results are reported as mean and standard deviation. We conduct experiments on a single NVIDIA A100 GPU.
@@ 2. $i$
Evaluation for different privacy-preserving VFL methods: Our method remains effective when applied to privacy-preserving VFL models. To validate this, we evaluate performance under Differential Privacy and Gradient Compression, two widely adopted mechanisms in VFL training (Fu et al., 2022a). The results confirm that our approach maintains robustness, underscoring its practicality for real-world deployments. Please refer to Appendix A.5 for a complete result.